solve:
5y-3x=14;3y-2x=1

Answer by mananth(16949)

5y-3x=14.............1
3y-2x=1.............2
..
multiply (1) by -2 & (2) by 3
-10y+6x=-28
9y-6x=3
add them up
-y =-25
y=25
...
plug value of y in (1)
5*25-3x=14
125-3x=14
-3x=-125+14
-3x=-111
/-3
x=37